A particle begins to move with a tangential acceleration of constant magnitude 0.6 m/s^2 in a circular path. If it slips when its total acceleration becomes 1 m/s. Find the angle through which it would have turned before it starts to slip.

Grade:12

1 Answers

HARSHDEEP SINGH CHHABRA
5 years ago
When net acceleration=1
Tangential acceleration=0.6 and centripetal =0.8
Now w^2R=0.8 
So w^2=0.8/r
 
Tagential acceleration=0.6=rα 
So α=0.6/r
Now as acceleration is constant, we can use
w^2=2α theta 
You can get theta from here which equals 2/3rad